# Share the Sugar

### v1

@@ 5 Incomparable outcomes
In the last section, we focused on incompleteness in the personal ranking of lives, and those forms of incompleteness in the impersonal ranking of outcomes that arise from it.
But incomparability between outcomes can have other sources as well. For example, those who resist interpersonal tradeoffs or aggregation might hold that, in some or all cases where $o$ is better than $o^{\prime}$ for some individuals and worse for others, $o$ and $o^{\prime}$ are impersonally incomparable. Or one might hold that there are imprecise “exchange rates” between different desirable features of a population, like total vs. average welfare or average welfare vs. equality, so that pairs of populations that trade off these dimensions are sometimes incomparable. (Critical-range utilitarianism, discussed above, can be seen as a view of this kind, in which “small population size” is a desideratum that trades off imprecisely with total welfare.)
This section will show that, as long as population-level incomparability has a plausible “separability” property analogous to Incomparability Transmission, these views are vulnerable to essentially the same difficulty that we explored in the last section. In other words, Negative Dominance and Personal Good support a general argument against incompleteness in impersonal value, whether or not it arises from incompleteness in personal value.
Our argument will roughly parallel the argument in the last section, with subpopulations taking the place of single individuals. A subpopulation is a set of possible individuals that have been arbitrarily indexed (i.e., assigned positive integers from 1 up to the number of individuals in the subpopulation). We also introduce the concept of a distribution, which stands in approximately the relation to subpopulations that lives stand in to individuals. A distribution is a function from the first $n$ positive integers to the set of lives (including nonexistence). The size of a subpopulation is the number of possible individuals it contains, and the size of a distribution is the number of lives it assigns. For any subpopulation $S$ consisting of possible individuals $S_{1},S_{2},\ldots,S_{n}$, and any distribution $D$ of size $n$ or less, we say that $S$ realizes $D$ when $S_{1}$ experiences the life $D(1)$, $S_{2}$ experiences the life $D(2)$, and so on, with any leftover individuals in $S$ being assigned nonexistence.
@@
Distributions $A$ and $B$ are strongly incomparable if, whenever outcomes $o$ and $o^{\prime}$ are identical except that some subpopulation realizes distribution $A$ in outcome $o$ and distribution $B$ in outcome $o^{\prime}$, $o$ is impersonally incomparable with $o^{\prime}$.
@@
It is plausible that, whenever two outcomes are incomparable, the two distributions they realize (relative to a given indexing of the set of individuals who exist in either outcome) are strongly incomparable. This would amount, however, to a fairly strong “separability” assumption, to the effect that in comparing two outcomes, we can focus only on those individuals whose lives differ between the outcomes and ignore everyone else. We choose, therefore, to focus on the weaker claim that, if impersonal betterness is incomplete, then at least some distributions are strongly incomparable. For example, perhaps a sufficiently sharp tradeoff between total welfare and inequality always generates incomparability (e.g., a choice between giving everyone in a particular subpopulation a modestly good life or giving half of them excellent lives and the other half mediocre lives), even if for less sharp tradeoffs, the impersonal betterness ranking can depend on features of the unaffected population. As we will see, just as in the last section, our puzzle does not require the full force of even this modest existential claim.
We now come to our result:
Proposition 5.1 (Population-Level Incomparability).
@@
Suppose that $A$ and $B$ are strongly incomparable. Letting $n$ be the size of the larger of $A$ and $B$, choose two disjoint subpopulations $S=\{S_{1},S_{2},\ldots,S_{n}\}$ and $T=\{T_{1},T_{2},\ldots,T_{n}\}$. Now, consider the two lotteries in Table 3.
@@
$T$ $B$ $A$
@@
$T$ $A$ $B$
@@
These are exact analogues of the two lotteries used to prove Proposition 4.1, with individuals $P_{1}$ and $P_{2}$ replaced by subpopulations $S$ and $T$, and lives $a$ and $b$ replaced by distributions $A$ and $B$. Each outcome assigns a distribution to each subpopulation: in the Heads outcome of Lottery 1, for instance, $S_{1}$ receives $A(1)$, $S_{2}$ receives $A(2)$, etc, and likewise $T_{1}$ receives $B(1)$, $T_{2}$ receives $B(2)$, etc, with any leftover individuals being assigned nonexistence.
Now, it follows from the definition of strong incomparability that $(A,B)$ and $(B,A)$ are both incomparable with $(A,A)$ and $(B,B)$ (where the first and second terms of each ordered pair represent the distributions realized by $S$ and $T$ respectively). By Small Improvements, there exists either an outcome $(A^{+},A)$ or $(A,A^{+})$ that slightly improves one life in $(A,A)$, or else an $(A^{-},A)$ or $(A,A^{-})$ that slightly worsens one life in $(A,A)$, that is still incomparable with both $(A,B)$ and $(B,A)$. As before, we focus on the first case. Richness guarantees that the domain of the impersonal betterness relation includes the two lotteries in Table 3. And in comparing these two lotteries, Negative Dominance implies that Lottery 2 is not better than Lottery 1, while Personal Good implies that it is.
As before, we note that all our argument really requires is that we can find distributions $A$, $A^{+}$, and $B$ such that $(A^{+},A)$ and $(B,B)$ are both incomparable with $(A,B)$ and $(B,A)$. We think it will be hard for proponents of incomparability between outcomes to deny this claim (we ourselves cannot think of any plausible incomplete outcome axiology that does so), even if they reject our Small Improvements assumption or do not believe that any distributions are strongly incomparable. It seems hard, therefore, to maintain that the impersonal betterness ranking of outcomes is incomplete without denying either Negative Dominance or Personal Good. [^20] Note also that the preceding argument applies to infinite distributions just as much as to finite distributions. Thus it rules out, for instance, strong incomparability between distributions like $(1,0,1,0,\ldots)$ and $(0,1,0,1,\ldots)$.
@@ 6 From outcomes to lotteries
In the last section, we presented an argument for completeness in the impersonal betterness ranking of outcomes. But the impersonal betterness relation, we have assumed, is defined not just on outcomes but on lotteries. And a complete ordering of outcomes need not entail a complete ordering of lotteries. For instance, even if outcomes are completely ordered by total welfare, the ordering of lotteries might be incomplete because it supervaluates over a range of reasonable risk attitudes. At an extreme, for instance, lotteries might be ordered only by Stochastic Dominance (Tarsney, 2020).
In this section, however, we take our argument a step further by showing that Personal Good goes a long way toward bridging the gap between completeness for outcomes and completeness for lotteries.Thus, if we accept completeness for outcomes on grounds that include Personal Good, there is at least significant reason to accept completeness for lotteries as well. This argument has somewhat limited scope (limited to what we will call “finite rational lotteries”), and it depends more heavily than the arguments of the last two sections on assumptions that go beyond Negative Dominance and Personal Good. So we do not think that this pair of principles support completeness for lotteries as strongly as they support completeness for lives and outcomes. But as we will see, the auxiliary assumptions we require are still quite plausible.
@@
For any outcome $o$ and any sufficiently large subpopulation $S$ (where what counts as “sufficiently large” can depend on $o$), there is an outcome exactly as impersonally good as $o$ in which only individuals in $S$ exist.
@@
If $o_{1}$ is at least as impersonally good as $o_{2}$, and $o_{2}$ is at least as impersonally good as $o_{3}$, then $o_{1}$ is at least as impersonally good as $o_{3}$.
Transitivity is a familiar principle that is very widely (though not universally) accepted, and we won’t have anything new to say for or against it. Substitution says, intuitively, that any degree of impersonal value can be realized by any sufficiently large group of possible individuals. This principle might be denied, for instance, by those who think that populations in which different individuals exist are always incomparable. But against the background assumption of completeness for outcomes, it is hard to see how one might deny it. Given completeness, in particular, Substitution follows from
@@
For any outcome $o$ and permutation $\pi$ of the set of possible individuals, $o$ is neither strictly better nor strictly worse than $\pi(o)$ (where $\pi(o)$ is the outcome in which $\pi(o)(x)=o(\pi(x))$.
Weak Anonymity seems like a minimal expression of impartiality. Given completeness for outcomes, Substitution can be understood as a still weaker sort of anonymity principle, which permits significant (though not unlimited) partiality between possible individuals.
Our argument will be restricted to finite rational lotteries: lotteries that have only finitely many possible outcomes, with each outcome having a rational probability. The “finitude” restriction is substantive, but we regard the “rationality” restriction as merely technical—it seems obvious that if lotteries with rational probabilities are completely ordered, then lotteries with real probabilities are as well.
@@
$s_{1}$ $\left(\frac{1}{n}\right)$ $s_{2}$ $\left(\frac{1}{n}\right)$ldots $s_{n}$
@@
The next step is to replace each outcome of each lottery with an equally good outcome, using “fresh” sets of individuals, so that no possible individual exists in more than one state in a given lottery. (It doesn’t matter whether the set of individuals who might exist in one lottery overlaps with the set of individuals who might exist in the other.) Call the resulting lotteries $L_{1^{\prime}}$ and $L_{2^{\prime}}$. These transformed lotteries are illustrated in Table 5. Substitution (together with Richness) guarantees that we can find lotteries with the required properties, since for each outcome of each lottery, we can find an equally good outcome involving a fresh set of individuals. Stochastic Dominance guarantees that $L_{1^{\prime}}$ is exactly as good as $L_{1}$, and $L_{2^{\prime}}$ is exactly as good as $L_{2}$.
@@
The final step is to “stack” all the individual outcomes other than nonexistence, in each lottery, into state $s_{1}$. Since all the possible outcomes of $L_{1^{\prime}}$ involve disjoint populations, and likewise for $L_{2^{\prime}}$, this stacking is possible. The result is a pair of lotteries $L_{1^{\prime\prime}}$ and $L_{2^{\prime\prime}}$ which guarantee an “empty” outcome in every state except $s_{1}$, while yielding a potentially massive population in that single state. These lotteries are illustrated in Table 6. Since all the states are equiprobable, this second transformation does not affect any individual’s personal lottery, so Personal Good implies that $L_{1^{\prime\prime}}$ is exactly as good as $L_{1^{\prime}}$, and likewise $L_{2^{\prime\prime}}$ is exactly as good as $L_{2^{\prime}}$.
@@
Let $o_{1}$ designate the outcome of $L_{1^{\prime\prime}}$ in state $s_{1}$, and $o_{2}$ designate the outcome of $L_{2^{\prime\prime}}$ in state $s_{1}$. Since $L_{1^{\prime\prime}}$ and $L_{2^{\prime\prime}}$ yield the same outcome in every state except $s_{1}$, Stochastic Dominance implies that, if $o_{1}$ is at least as good as $o_{2}$, then $L_{1^{\prime\prime}}$ is at least as good as $L_{2^{\prime\prime}}$; and likewise, if $o_{2}$ is at least as good as $o_{1}$, then $L_{2^{\prime\prime}}$ is at least as good as $L_{1^{\prime\prime}}$. Since we have assumed that outcomes are completely ordered, we know that at least one of these conditions obtains. Therefore, we know that either $L_{1^{\prime\prime}}$ is at least as good as $L_{2^{\prime\prime}}$ or vice versa (or both). Finally, Transitivity implies that $L_{1}$ is exactly as good as $L_{1^{\prime\prime}}$, $L_{2}$ is exactly as good as $L_{2^{\prime\prime}}$, and therefore that $L_{1}$ is at least as good as $L_{2}$ if and only if $L_{1^{\prime\prime}}$ is at least as good as $L_{2^{\prime\prime}}$. Thus either $L_{1}$ is at least as good as $L_{2}$, or $L_{2}$ is at least as good as $L_{1}$, or both. Since $L_{1}$ and $L_{2}$ were chosen arbitrarily, this shows that finite rational lotteries are completely ordered.
@@
If you were convinced by the arguments for outcome completeness in the last two sections, should you be convinced by this argument as well? We think so. The additional premises are Stochastic Dominance, Substitution, and Transitivity. Stochastic Dominance is, as we have said, a very plausible and widely accepted principle. And, although it neither implies nor is implied by Negative Dominance, both principles reflect the basic idea that we should evaluate lotteries based on the overall value and probability of their possible outcomes, so the principles have some overlapping appeal. [^22] On the other hand, those who accept Negative Dominance because they accept the stronger principle of Statewise Negative Dominance will, as we have seen, reject Stochastic Dominance. But we don’t think they’ll reject the particular uses we make of it in our proof. In fact, in the framework of states and gambles, we could replace Stochastic Dominance in the preceding argument with Strong Statewise Dominance If in every state, the outcome of gamble $G_{1}$ is at least as good as the outcome of gamble $G_{2}$, then $G_{1}$ is at least as good as $G_{2}$. If in addition there is some state in which the outcome of $G_{1}$ is strictly better than the outcome of $G_{2}$, then $G_{1}$ is strictly better than $G_{2}$. Anyone who accepts Statewise Negative Dominance should, we think, find this principle compelling. (Our argument would then also require the innocuous principle that subdividing a state, without changing the outcome of a given gamble in that state, does not affect the value of the gamble.) Transitivity is likewise widely accepted. And although Substitution could certainly be questioned, it is hard to reject if we have already accepted completeness for outcomes, since it is then entailed by Weak Anonymity.
7 Responses
The last three sections have shown that, against very weak background assumptions, Negative Dominance and Personal Good, are incompatible with various forms of value incompleteness. So far, we have presented our results as arguments against these forms of incompleteness, and we expect that many readers will wish to take them this way. But, of course, one could also choose to give up Negative Dominance or Personal Good.In this section we will consider both possibilities, focusing especially on the former.
There are at least three reasons, beyond intuitions about cases, why one might be motivated to hang onto incomparability despite the preceding arguments. First, it seems strange that principles about the betterness ordering of lotteries (like Negative Dominance and Personal Good) should constrain the betterness ordering of outcomes. This seems backwards, given the plausible thought that the betterness of outcomes is more fundamental (with betterness of lotteries being grounded in betterness of outcomes plus decision-theoretic principles of rational preference under risk). It would seem particularly odd if principles about the impersonal betterness of lotteries forced us to accept particular conclusions about the personal betterness of lives—the comparative personal value of, say, the lives of an artist and a banker seems like a prior question to principles like Negative Dominance and Personal Good. If those lives are incomparable, one might think, that is a fact that ethical and decision-theoretic principles must bend to accommodate.
Second, there are powerful theoretical arguments for incompleteness in certain infinitary contexts: “constructive incompleteness” results in infinite ethics (Zame, 2007; Lauwers, 2010) and the decision theory of lotteries without finite expected value (Lauwers, 2016) show that given some relatively weak assumptions, any ranking of infinite outcomes or lotteries respectively must be either incomplete or non-constructive (having no finite description). Many find constructiveness a very compelling constraint on a theory of value: a non-constructive betterness relation seems objectionably arbitrary (for instance, in the context of infinite ethics, it would require infinitely many apparently-arbitrary choices about which of two countably infinite sets of possible individuals should be prioritized over the other) and would be impossible for any finite agent to represent, let alone know.
Third and finally, even without incompleteness, Personal Good and Negative Dominance conflict with one another in infinitary contexts—-when comparing lotteries with infinitely many possible outcomes and that involve infinite or unboundedly large populations. For example, Table 7 describes a pair of lotteries in which, for every possible individual, the personal lottery given by Lottery 1 stochastically dominates the personal lottery given by Lottery 2; and yet Lottery 2 is certain to result in a positive total of personal value, while Lottery 1 is certain to result in a negative total of personal value. Thus Personal Good implies that Lottery 1 is better than Lottery 2 while, at least given a total utilitarian theory of impersonal value, Negative Dominance implies that it cannot be. [^23] This example is inspired by one inKowalczyk . In Kowalczyk’s example, Personal Good conflicts with Statewise Equivalence (the principle that, if two gambles have equally good outcomes in every state, then they’re equally good). We have modified the example to create a conflict with Negative Dominance. (Indeed, in our case, Personal Good conflicts with every dominance principle we can think of. In particular, it conflicts with Superdominance—the principle that, if each possible outcome of $L_{i}$ is strictly better than every possible outcome of $L_{j}$, then $L_{i}$ is better than $L_{j}$. A fortiori, it conflicts with Stochastic Dominance and Statewise Dominance as well.) Related examples of conflicts between dominance principles and ex ante Pareto-like principles appear in Goodsell (2021), Wilkinson (2023), and Nebel (2023). While we have used totalism to illustrate the argument in the main text, the axiological assumptions required to generate this sort of conflict between Personal Good and Negative Dominance are very weak. It would suffice, for instance, to assume a principle we might call “Inverse Lives”: There exists a pair of lives $l^{+}$ and $l^{-}$ such that, among outcomes in which everyone who exists has one of those two lives, any outcome where most people have $l^{+}$ is better than any outcome where most people have $l^{-}$. This assumption is endorsed, for instance, by standard versions of total utilitarianism, critical-level utilitarianism, prioritarianism, average utilitarianism, variable-value utilitarianism, and many forms of egalitarianism. Finally, we have presented the example as involving only finite (though unboundedly large) populations. We could switch to an infinite-population example by replacing every instance of nonexistence with Table 7 with a neutral life, with welfare 00. This would allow us to weaken Personal Good by restricting it to fixed-population contexts, where every individual who might exist in either lottery is certain to exist in both lotteries. Then the principle could simply assert that if one lottery gives every individual a stochastically dominant personal lottery, then it’s impersonally better. While it is not at all obvious which principle we should give up in these contexts, there is strong reason to think that Personal Good and Negative Dominance cannot both be completely universal truths. Whichever principle we choose to give up in infinite contexts, this should arguably undercut its intuitive appeal in finite contexts as well, and make us more willing to sacrifice that principle in order to preserve space for incomparability.

### v2

@@ 5 Incomparable outcomes
In the last section, we focused on incompleteness in the personal ranking of lives, and those forms of incompleteness in the impersonal ranking of outcomes that arise from it. But incomparability between outcomes can have other sources as well. For example, those who resist interpersonal tradeoffs or aggregation might hold that, in some or all cases where $o_{1}$ is better than $o_{2}$ for some individuals and worse for others, $o_{1}$ and $o_{2}$ are incomparable. Or one might hold that there are imprecise “exchange rates” between different desirable features of an outcome, like total vs. average welfare or average welfare vs. equality, so that pairs of outcomes that trade off these dimensions are sometimes incomparable. (Critical-range utilitarianism, discussed above, can be seen as a view of this kind, in which “small population size” is a desideratum that trades off imprecisely with total welfare.)
This section will show that, as long as population-level incomparability has a plausible “separability” property analogous to Incomparability Transmission, these views are vulnerable to essentially the same difficulty that we explored in the last section. In other words, Negative Dominance and Personal Good support a general argument against incompleteness in the impersonal value of outcomes, whether or not it arises from incompleteness in personal value.
Our argument will roughly parallel the argument in the last section, with subpopulations taking the place of single individuals. A subpopulation is a set of possible individuals that have been arbitrarily indexed (i.e., assigned positive integers from 1 up to the number of individuals in the subpopulation). We also introduce the concept of a distribution, which stands in approximately the relation to subpopulations that lives stand in to individuals. A distribution is a function from the first $n$ positive integers to the set of lives (including nonexistence). The size of a subpopulation is the number of possible individuals it contains, and the size of a distribution is the number of lives it assigns. For any subpopulation $S$ consisting of possible individuals $P_{S1},P_{S2},\ldots,P_{Sn}$, and any distribution $D$ of size $n$ or less, we say that $S$ realizes $D$ when $P_{S1}$ experiences the life $D(1)$, $P_{S2}$ experiences the life $D(2)$, and so on, with any leftover individuals in $S$ being assigned nonexistence.
@@
Distributions $A$ and $B$ are strongly incomparable if and only if, whenever outcomes $o_{1}$ and $o_{2}$ are identical except that some subpopulation realizes distribution $A$ in outcome $o_{1}$ and distribution $B$ in outcome $o_{2}$, $o_{1}$ is incomparable with $o_{2}$.
@@
It seems plausible that, whenever two outcomes are incomparable, the two distributions they realize (relative to a given indexing of the set of individuals who exist in either outcome) are strongly incomparable. But this is a substantial assumption, amounting to a form of separability: it allows us, in comparing two outcomes, to focus only on those individuals whose lives differ between the outcomes and ignore everyone else. Since this assumption may be controversial, we will focus on the much weaker claim that, if impersonal betterness is incomplete, then at least some distributions are strongly incomparable. For example, perhaps a sufficiently sharp tradeoff between total welfare and inequality in a subpopulation of a given size always generates incomparability, even if for less sharp tradeoffs, the impersonal betterness ranking can depend on features of the unaffected population. As we will see, our argument will not require the full force of even this modest existential claim.
We now come to our next result:
Proposition 3.
@@
Suppose that $A$ and $B$ are strongly incomparable. Letting $n$ be the size of the larger of $A$ and $B$, choose two disjoint subpopulations $S=\{P_{S1},P_{S2},\ldots,P_{Sn}\}$ and $T=\{P_{T1},P_{T2},\ldots,P_{Tn}\}$. Now, consider the two lotteries in Table 5.
@@
$T$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}A}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B}$
@@
$T$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}A}$
@@
These are exact analogues of the two lotteries used to prove Proposition 1, with individuals $P_{1}$ and $P_{2}$ replaced by subpopulations $S$ and $T$, and lives $a$ and $b$ replaced by distributions $A$ and $B$. Each outcome assigns a distribution to each subpopulation: in the Tails outcome of Lottery 2, for instance, $P_{S1}$ receives $B(1)$, $P_{S2}$ receives $B(2)$, etc, and likewise $P_{T1}$ receives $A(1)$, $P_{T2}$ receives $A(2)$, etc, with any leftover individuals being assigned nonexistence.
Now, it follows from the definition of strong incomparability that $(A,A)$ and $(B,B)$ are both incomparable with $(A,B)$ and $(B,A)$ (where the first and second terms of each ordered pair represent the distributions realized by $S$ and $T$ respectively). By Small Improvements, there exists either an outcome $(A^{+},B)$ or $(A,B^{+})$ that slightly improves one life in $(A,B)$, or else an $(A^{-},B)$ or $(A,B^{-})$ that slightly worsens one life in $(A,B)$, that is still incomparable with both $(A,A)$ and $(B,B)$. As before, we focus on the first case. Richness guarantees that the domain of the impersonal betterness relation includes the two lotteries in Table 5. And in comparing these two lotteries, Negative Dominance implies that Lottery 2 is not better than Lottery 1, while Personal Good implies that it is.
As before, all our argument really requires is that we can find distributions $A$, $A^{+}$, and $B$ such that $(A,A)$ and $(B,B)$ are both incomparable with $(A^{+},B)$ and $(B,A)$. We think it will be hard for proponents of incomparability between outcomes to deny this claim (we ourselves cannot think of any plausible incomplete outcome axiology that does so), even if they reject our Small Improvements assumption or deny that any distributions are strongly incomparable. It seems hard, therefore, to maintain that the impersonal betterness ranking of outcomes is incomplete without denying either Negative Dominance or Personal Good.
@@ 6 From outcomes to lotteries
In the last section, we presented an argument for completeness in the impersonal betterness ranking of outcomes. But the impersonal betterness relation, we have assumed, is defined not just on outcomes but on lotteries. And a complete ordering of outcomes need not entail a complete ordering of lotteries. For instance, even if outcomes are completely ordered by total welfare, the ordering of lotteries might be incomplete because it supervaluates over a range of reasonable risk attitudes. At an extreme, for instance, lotteries might be ordered only by Stochastic Dominance (Tarsney 2020).
In this section, we take our argument a step further by showing that Personal Good goes a long way toward bridging the gap between completeness for outcomes and completeness for lotteries. Thus, if we accept completeness for outcomes on grounds that include Personal Good, there is at least significant reason to accept completeness for lotteries as well. This argument has somewhat limited scope (limited to what we will call “finite rational lotteries”), and it depends more heavily than the arguments of the last two sections on assumptions that go beyond Negative Dominance and Personal Good. So we do not think that this pair of principles support completeness for lotteries as strongly as they support completeness for lives and outcomes. But as we will see, the auxiliary assumptions we require are still quite plausible.
@@
For any outcome $o$ and any sufficiently large subpopulation $S$ (where what counts as “sufficiently large” can depend on $o$, but is always finite) there is an outcome that is equally as good as $o$ in which only individuals in $S$ exist.
@@
If $L_{1}$ is at least as good as $L_{2}$, and $L_{2}$ is at least as good as $L_{3}$, then $L_{1}$ is at least as good as $L_{3}$.
Transitivity is a familiar principle that is very widely (though not universally) accepted, and we won’t have anything new to say for or against it. Substitution says, intuitively, that any degree of impersonal value can be realized by any sufficiently large group of possible individuals. This principle might be denied, for instance, by those who think that populations in which different individuals exist are always incomparable. But against the background assumption of completeness for outcomes, which we have just argued for, it’s hard to see how one might deny it. Given completeness, in particular, Substitution follows from
@@
For any outcome $o$ and permutation $\pi$ of the set of possible individuals, $o$ is neither better nor worse than $\pi(o)$ (where $\pi(o)$ is the outcome in which $\pi(o)(x)=o(\pi(x))$.
Weak Anonymity seems like a minimal expression of impartiality. Given completeness for outcomes, Substitution can be understood as a still weaker sort of anonymity principle, which permits significant (though not unlimited) partiality between possible individuals.252525 The background restriction to finite-population outcomes is important here because, as with other anonymity principles, Substitution becomes much stronger and easier to deny in infinite-population contexts. For instance, consider an outcome $o$ in which a countably infinite set $T$ of possible individuals exist, all with lives that are maximally or near-maximally good (assuming that to be possible). And consider a subset $S$ of $T$ such that $S$ and its relative complement in $T$ are both infinite. Although $S$ is as large, in cardinal terms, as a subpopulation can get (given the assumption of only countably many possible individuals), it is at least doubtful whether any outcome in which only members of $S$ exist can be as good as $o$.
Our argument will be restricted to finite rational lotteries: lotteries over finite-population outcomes that are finitely supported (i.e., have only finitely many possible outcomes), with each outcome having a rational probability. The finitary restrictions are substantive, and we make them explicit here (although they have been implicitly in force throughout) for two reasons. First, this is the only one of our results that formally depends on finitary restrictions—in particular, the proof below requires the assumption of finite support. Second, while the result could formally be extended to finite-support lotteries over infinite populations, there is particular reason here to be skeptical of our axioms in that context—see n. 25. By contrast, we regard the “rationality” restriction as merely technical—it seems obvious that if lotteries with rational probabilities are completely ordered, then lotteries with real probabilities are as well.
@@
The next step is to replace each outcome of each lottery with an equally good outcome, using “fresh” sets of individuals, so that no possible individual exists in more than one outcome of a given lottery. (It doesn’t matter whether the set of individuals who might exist in one lottery overlaps with the set of individuals who might exist in the other.) Call the resulting lotteries $L_{1^{\prime}}$ and $L_{2^{\prime}}$. These transformed lotteries are illustrated in Table 7. Substitution (together with Richness) guarantees that we can find lotteries with the required properties, since for each outcome of each lottery, we can find an equally good outcome involving a fresh set of individuals. Stochastic Dominance guarantees that $L_{1^{\prime}}$ and $L_{1}$ are equally good, and that $L_{2^{\prime}}$ and $L_{2}$ are equally good.
@@
The final step is to “stack” all the individual outcomes other than nonexistence, in each lottery, into state $s_{1}$. Since all the possible outcomes of $L_{1^{\prime}}$ involve disjoint populations, and likewise for $L_{2^{\prime}}$, this stacking is possible. The result is a pair of lotteries $L_{1^{\prime\prime}}$ and $L_{2^{\prime\prime}}$ which guarantee an “empty” outcome in every state except $s_{1}$, while yielding a potentially massive population in that single state. These lotteries are illustrated in Table 8. Since all the states are equiprobable, this second transformation does not affect any individual’s personal lottery, so Personal Good implies that $L_{1^{\prime\prime}}$ and $L_{1^{\prime}}$ are equally good, and likewise that $L_{2^{\prime\prime}}$ and $L_{2^{\prime}}$ are equally good.
@@
Let $o_{1}$ designate the outcome of $L_{1^{\prime\prime}}$ in state $s_{1}$, and $o_{2}$ designate the outcome of $L_{2^{\prime\prime}}$ in state $s_{1}$. Since $L_{1^{\prime\prime}}$ and $L_{2^{\prime\prime}}$ yield the same outcome in every state except $s_{1}$, Stochastic Dominance implies that, if $o_{1}$ is at least as good as $o_{2}$, then $L_{1^{\prime\prime}}$ is at least as good as $L_{2^{\prime\prime}}$; and likewise, if $o_{2}$ is at least as good as $o_{1}$, then $L_{2^{\prime\prime}}$ is at least as good as $L_{1^{\prime\prime}}$. Since we have assumed that outcomes are completely ordered, we know that at least one of these conditions obtains. Therefore, we know that either $L_{1^{\prime\prime}}$ is at least as good as $L_{2^{\prime\prime}}$ or vice versa (or both). Finally, Transitivity implies that $L_{1}$ and $L_{1^{\prime\prime}}$ are equally good, $L_{2}$ and $L_{2^{\prime\prime}}$ are equally good, and therefore that $L_{1}$ is at least as good as $L_{2}$ if and only if $L_{1^{\prime\prime}}$ is at least as good as $L_{2^{\prime\prime}}$. Thus either $L_{1}$ is at least as good as $L_{2}$, or $L_{2}$ is at least as good as $L_{1}$, or both. Since $L_{1}$ and $L_{2}$ were chosen arbitrarily, this shows that finite rational lotteries are completely ordered.
@@
Should a reader convinced by the arguments for outcome completeness in the last two sections be convinced by this argument as well? We think so. The additional premises are Stochastic Dominance, Substitution, and Transitivity. Stochastic Dominance is, as we have said, a very plausible and widely accepted principle. And, although it neither implies nor is implied by Negative Dominance, both principles reflect the basic idea that we should evaluate lotteries based on the values and probabilities of their possible outcomes, so the principles have some overlapping appeal. Transitivity is likewise widely accepted. And although Substitution could certainly be questioned, it is hard to reject if we have already accepted completeness for outcomes, since it is then entailed by Weak Anonymity, which is widely accepted for finite populations.
7 Objections
The last three sections have shown that, against weak background assumptions, Negative Dominance and Personal Good are incompatible with incompleteness in the personal value of lives and the impersonal value of finite outcomes and lotteries. We see this as a compelling argument against those forms of incompleteness. But of course defenders of incompleteness might choose to reject Negative Dominance or Personal Good instead.  [^27] Indeed, we should note that, while the authors of this paper all see Negative Dominance and Personal Good as compelling principles, we ourselves are not uniformly certain of their correctness. In particular, one of us (CT) has some idiosyncratic reservations about Personal Good.
In this section we will consider three objections to our argument that might motivate rejecting one of these principles rather than incompleteness, and explain why we find them unconvincing.
Objection
An intuitively strange feature of the preceding arguments is that they use principles about the ranking of lotteries (Negative Dominance and Personal Good) to derive strong constraints on the rankings of lives and outcomes. This might seem to get things backwards: The betterness relations on lives and outcomes are more fundamental than the betterness relation on lotteries and so, one might claim, principles about the latter relation should be constrained by rather than constrain the former relations. Moreover, personal betterness is more fundamental than impersonal betterness, so it’s especially strange for principles about the impersonal ranking of lotteries to constrain the personal ranking of lives. The comparative personal value of, say, the lives of an artist and a banker is a prior question to principles like Negative Dominance and Personal Good. If those lives are incomparable, that is a fact that ethical and decision-theoretic principles must bend to accommodate, not something that should be denied because it forces ethicists and decision theorists to make difficult choices.
Response
This argument rests on an implicit premise we reject: that the relevant kind of fundamentality—be it metaphysical or some other form of ethical fundamentality—restricts what sorts of propositions can be used as evidence for which others. Our argument does not rely on the claim that, metaphysically speaking, facts about the betterness of lotteries ground or explain the betterness of lives and outcomes. It requires only that structural principles governing the ranking of lotteries can give us evidence about structural principles governing the rankings of lives and outcomes. Elementary particles are metaphysically more fundamental than galaxies, but we can still revise our beliefs about elementary particles on the basis of evidence provided by the behavior of galaxies.
Objection
The arguments of §§45 relied on the combination of Negative Dominance and Personal Good. Although these two principles can be jointly accepted in the finite context that has been assumed thus far, they turn out to be inconsistent in certain infinitary contexts—specifically, when comparing lotteries with infinitely many possible outcomes that involve infinite or unboundedly large populations.